The accounting department of your company has just delivered a drat of the current years fnancial statements to you The summary is as folows You discovered that they have not adjusted for estimated bad debt expenses of $8.000 For each of the tollowing ratios, cakculiate 1. The ratio that would have resulted had the error not been discovered (.e. the incorlect rato) 2. The correct ratio 2 ROA 3 ROE 3631% 464 36 84% 5 EPS
Answer #1

Ans:

Particulars

Incorrect

Correct

ROA

Net Income/average assets = 97,500/(550000+559000/2) = 97,500/554500 = 17.58%

=(97,500-8000) / [(550000+559000-8000)/2] = 89,500/550,500 = 16.26%

ROE

Net Income/average equity = 97,500/(340000+356,000/2) = 97,500/348,000 = 28.02

= (97,500-8000) / [(340000+356000-8000)/2] = 89500/344,000 = 26.02

Debt Ratio

Total debt / Average Assets = 203000 /559000) = 203000/559,000 = 36.31

=203000 / [(559000-8000)] = 203000/551000 = 36.84

EPS

Net Income / Number of Common Shares = 97,500/21000 = 97,500/20000 = $4.64/share

= (97,5000-8000)/21000 =89500/21000=$4.26/Share